Z80ASM(1L) Z80ASM(1L) NAME as64180,asz80,as8085,as8080−crossassemblersformicrocomputers SYNOPSIS as64180[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input asz80[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input as8085[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input as8080[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input DESCRIPTION Theas64180,asz80,as8085,as8080command(s)assemblestheinputfileintoatextoutputfile representingtheprogrammemoryforamicrocomputer. Options -pcpuOverridetheinstructionsetselection. ValidValuesforCpu 64180,z80,Z80,8085,8085 -llistfile Outputafileformatedwiththememoryaddress,data,andsourceinputlines. -ohexfile Outputthememorydataimageinaformacceptedbymostpromprogrammers. -hhexfile Sameas−o. −dSavetheintermediatedatafile(seeFILE)andaborttheexecutionattheendofprocessing. -ssymbolfile Printthesymboltablevaluesandnames,oneperlineinthespecifiedfile. FILES /usr/tmp/frtXXXXXX SEEALSO FrankensteinCrossAssemblersUsersManual(filebase.doc)Appendixforasz80Frankenstein Assembler(fileasz80.doc) NOTES Thereisonlyoneinputfile. Theprogramcanselectwhichsubsetofinstructionsisaccepted.Theprogramfirstusesthename withwhichitisinvoked.Thisisscannedforasubstringwhichindicateswhichsettouse.The-p optionsoverridesthisselctionbyperformingthesamesubstringsearch.Finallytheinputfilecan selectwhichsubsettousewiththeCPUstatement. Thereshouldonlybeoneexecutablefilewithallofitsnameslinkedtoit. 1